OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx +c=0 The solutions are -10, 4 The equation is __ =0

OpenStudy (anonymous):

\[ax^2 + bx +c=0\]

OpenStudy (anonymous):

get the sum of the roots, and product of the roots. -10+4 = -6 -10*4 = -40 So the answer is: x^2+6x-40 = 0

OpenStudy (anonymous):

if you have a quadratic equation in the form: \[x^2+bx+c = 0\] sum of the roots is -b, product of the roots is c